# Game-Theoretic Discovery of Quantum Error-Correcting Codes Through Nash Equilibria

### v2

@@ Abstract
Scalability to hardware-relevant system sizes is demonstrated at $n=100$ qubits—matching experimental scales of recent demonstrations [31]—where the framework discovers codes across all six objectives with tractable computational cost. Distance optimization discovers [[100,49,3]][\![100,49,3]\!] codes with 49% encoding rate. Hardware adaptation produces [[100,50,3]][\![100,50,3]\!] codes respecting 2D connectivity constraints with maximum degree $\Delta\leq 3$. Rate-distance tradeoff and cluster-state search achieve [[100,50,4]][\![100,50,4]\!] codes with distance-4 error protection and 50% encoding rate, significantly surpassing the [[15,7,3]][\![15,7,3]\!] validation example in both scale and distance. Surface-like topology optimization generates [[100,46,4]][\![100,46,4]\!] codes with average degree near 4. Discovery at $n=100$ completes in 40–66 minutes for 20 independent trials with 50 iterations each using 28 parallel workers on standard workstations, with convergence typically occurring within 20–40 iterations. Computational cost per iteration scales as $O(n^{3})$, dominated by stabilizer rank computation via Gaussian elimination over $\mathbb{F}_{2}$, enabling practical code discovery at experimentally relevant scales versus exhaustive search complexity $O(2^{n^{2}})$ requiring $10^{13}$ seconds for $n=20$. These results demonstrate the framework operates efficiently at scales required for near-term fault-tolerant quantum computing demonstrations, complementing asymptotic QLDPC constructions by providing systematic exploration at finite length scales with mechanistic interpretability.
@@
Practical performance assessment employed syndrome-based error correction simulation for discovered codes against surface code baselines at hardware-relevant scales (Fig.˜3). Using belief propagation with ordered statistics decoding (BP+OSD) [32], we measured logical error rates $\varepsilon_{L}$ versus physical error rates $p$ for depolarizing noise $\varepsilon=(p/3)(X+Y+Z)$ applied independently to each qubit. Monte Carlo simulation with $10^{5}$ syndrome measurement cycles per data point (ensuring statistical uncertainties $\Delta\varepsilon_{L}/\varepsilon_{L}<5\%$) demonstrates that discovered codes achieve comparable error suppression to surface codes. Distance-3 discovered codes reach $\varepsilon_{L}=10^{-3}$ at $p=10^{-2}$, matching surface code performance (Fig.˜3 A). Distance-5 discovered codes achieve $\varepsilon_{L}=10^{-4}$ at $p=10^{-2}$, with threshold behavior $\varepsilon_{L}\propto p^{(d+1)/2}$ confirming distance scaling. The rediscovered [[15,7,3]][\![15,7,3]\!] code achieves 47% encoding rate ($k/n=7/15$), substantially higher than distance-3 surface codes encoding one logical qubit per $\sim$9 physical qubits ($k/n\approx 0.11$), providing resource efficiency advantages in overhead-constrained regimes. For hardware-adapted codes with degree $\Delta\leq 3$ on IBM heavy-hex topology [33], the [[15,7,3]][\![15,7,3]\!] code maintains $\varepsilon_{L}<10^{-3}$ at $p=10^{-3}$, meeting fault-tolerance thresholds while fitting existing hardware connectivity—a capability algebraic constructions rarely achieve simultaneously. Computational efficiency enables code discovery for $n\leq 20$ in under 1 second on standard workstations versus exhaustive search’s $O(2^{n^{2}})$ complexity requiring $10^{13}$ seconds for $n=20$ (Fig.˜3 C), with the tractable region extending to $n\approx 100$ through optimized implementations.
Figure 3: Practical Advantages at Hardware Scales. (A) Logical error rate $\varepsilon_{L}$ versus physical error rate $p$ for discovered codes (solid lines with circles/diamonds) and surface code baselines (dashed lines with squares) at distances $d=3,5$. Error bars show $\pm 1\sigma$ statistical uncertainty from Monte Carlo simulation with $10^{5}$ syndrome measurement cycles. Gray dashed line shows uncorrected error rate for reference. Discovered codes achieve comparable error suppression with threshold behavior $\varepsilon_{L}\propto p^{(d+1)/2}$. (B) Encoding rate comparison: the rediscovered [[15,7,3]][\![15,7,3]\!] quantum Hamming code achieves $k/n=0.47$ (47% encoding rate), substantially higher than distance-3 surface codes at $k/n\approx 0.11$, providing resource efficiency in overhead-constrained regimes. (C) Discovery time scaling with system size $n$. Game-theoretic approach (blue solid, $O(n^{3})$ per iteration) remains tractable through $n\approx 100$ (green-shaded region), crossing exhaustive search complexity (red dashed, $O(2^{n^{2}})$) at $n=7$ with over 6 orders of magnitude advantage at $n=20$.
Framework extensibility distinguishes it from specialized algorithms. Adding objectives requires only defining evaluation functions $f_{\text{new}}(G)$—no algorithmic restructuring. For example, introducing fault-tolerance objectives targeting transversal gate implementations [34] simply adds penalty terms for non-CSS structure or rewards for logical operator support overlap. Biased noise objectives $f_{\text{bias}}=d_{Z}^{2}\cdot(1+d_{X}/d_{Z})$ targeting asymmetric error protection for biased noise channels [35] immediately enable tailored code discovery by separately evaluating $X$- and $Z$-distance. Hardware topologies (2D grid, heavy-hex lattices, Rydberg atom arrays, all-to-all connectivity) constrain strategy spaces through graph planarity or degree restrictions without modifying core game logic. This modularity contrasts sharply with algebraic methods, where each code family (CSS, topological, LDPC) requires distinct mathematical machinery (classical codes, lattice geometry, expander graphs), and with reinforcement learning, where reward shaping, neural architecture, and training protocols must be redesigned for new objectives [13]. The transparent circuit construction enabled by graph states provides critical verification advantage. Unlike abstract stabilizer formalism requiring circuit compilation, the explicit mapping $G\rightarrow|G\rangle=\prod_{(u,v)\in E}\text{CZ}_{uv}|+\rangle^{\otimes n}$ allows immediate implementation. For the [[15,7,3]][\![15,7,3]\!] MBQC code, preparation requires 15 Hadamard gates followed by 21 CZ gates (one per edge), with syndrome measurement using 15 ancillas and circuit depth $\propto\Delta(G)=3$. This transparency facilitates experimental realization and independent verification—a significant advantage when proposing novel codes for near-term devices.
Our approach shares conceptual grounding with optimization via games [36] but applies game dynamics in reverse direction: rather than using quantum mechanics to play games [17, 18], we use game theory to construct quantum systems. Recent quantum game theory work focuses on quantum strategies providing computational advantages in game-playing scenarios, orthogonal to our construction methodology. The “quantum lights out” framework [20] employs single-player optimization metaphors but lacks strategic interactions between competing objectives and equilibrium concepts central to game theory. Reinforcement learning for code discovery [12, 14] explicitly avoids game-theoretic formulation, treating optimization as single-agent learning against deterministic environments: “Unlike chess where an agent plays against an opponent, the code game is deterministic.” Adversarial noise models in recent QEC literature [21] analyze robustness against worst-case errors but do not employ game theory for code construction. To our knowledge, no prior work bridges multi-agent game theory and quantum error correction for systematic code discovery through Nash equilibria.
@@
Several limitations warrant acknowledgment. Graph state restriction excludes important non-graph-state codes (five-qubit code, Steane code, Bacon-Shor codes) not representable with this formalism. Distance estimation employs heuristics based on stabilizer rank over $\mathbb{F}_{2}$ and graph connectivity that may underestimate true distance; we validate critical cases with exact weight enumeration when tractable. The framework currently optimizes code structure but not decoders; joint code-decoder optimization through co-evolutionary game dynamics represents future extension. Approximate Nash equilibria guarantee only local optimality; finding globally optimal codes would require exhaustive equilibrium search over all possible objective configurations, computationally prohibitive. The $n=100$ codes discovered, while demonstrating scalability, require verification against comprehensive code databases to determine novelty versus known constructions. Despite these constraints, the framework’s systematic rediscovery of the optimal [[15,7,3]][\![15,7,3]\!] quantum Hamming code validates the methodology, while scalability to $n=100$ demonstrates practical applicability at hardware-relevant scales, positioning the approach for discovering competitive codes in unexplored parameter regimes.
Extensions include decoder co-optimization through additional players representing decoder strategies, dynamic objective reconfiguration for degraded hardware where qubit failures trigger real-time game adjustments, and applications beyond error correction to quantum circuit optimization (gate count minimization via circuit topology games) and resource state design (graph state preparation for distributed quantum computing). The interpretable nature of equilibrium-driven discovery may accelerate transition from theoretical code families to practical implementations in near-term devices by providing mechanistic understanding of why certain topologies succeed—enabling physicists to refine codes based on experimental constraints rather than treating optimization as black-box search. This work opens research avenues at the intersection of game theory and quantum information, demonstrating that strategic interactions provide systematic, interpretable frameworks for quantum system design beyond traditional algebraic or computational approaches.

### v3

@@ D. Objective design principles
Scalability to hardware-relevant system sizes is demonstrated at $n=100$ qubits—matching experimental scales of recent demonstrations [30]—where the framework discovers codes across all seven objectives with tractable computational cost. Distance optimization discovers [[100,49,3]][\![100,49,3]\!] codes with 49% encoding rate. Hardware adaptation produces [[100,50,3]][\![100,50,3]\!] codes respecting 2D connectivity constraints with maximum degree $\Delta\leq 3$. Rate-distance tradeoff and cluster-state search achieve [[100,50,4]][\![100,50,4]\!] codes with distance-4 error protection and 50% encoding rate, significantly surpassing the [[15,7,3]][\![15,7,3]\!] validation example in both scale and distance. Surface-like topology optimization generates [[100,46,4]][\![100,46,4]\!] codes with average degree near 4. Discovery at $n=100$ completes in 40–66 minutes for 20 independent trials with 50 iterations each using 28 parallel workers on the workstation described above, with convergence typically occurring within 20–40 iterations. The $O(n^{3})$ per-iteration cost established above enables practical code discovery at experimentally relevant scales, versus exhaustive search complexity $O(2^{n^{2}})$ requiring $10^{13}$ seconds for $n=20$. These results demonstrate the framework operates efficiently at scales required for near-term fault-tolerant quantum computing demonstrations, complementing asymptotic QLDPC constructions by providing systematic exploration at finite length scales with mechanistic interpretability.
The preceding results establish that the framework discovers codes with favorable parameters across diverse objectives and system sizes. We now assess whether these codes provide practical value: metrological sensitivity quantified by $\mathcal{F}_{Q}$, and error suppression under realistic noise.
@@
Practical performance assessment combines metrological benchmarking with syndrome-based error correction simulation (Fig.˜5). Quantum Fisher information $\mathcal{F}_{Q}$ evaluated for Nash-discovered codes at $n=10$, 12, and 15 qubits reveals SQL scaling $\mathcal{F}_{Q}\approx n$ with power-law exponent $\alpha=1.00$ (Fig.˜5A), consistent with the distance-3 codes produced by the Fisher Information objective. All three system sizes yield $d=3$ codes that saturate the SQL, confirming that the game-theoretic framework identifies metrologically useful states across the range n=10n=10–15 without manual parameter tuning. Beyond the SDP-tractable regime ($n\leq 10$), the Nash approach provides, to our knowledge, the only systematic method for constructing codes with certified metrological performance. Using belief propagation with ordered statistics decoding (BP+OSD) [31], we measured logical error rates $\varepsilon_{L}$ versus physical error rates $p$ for depolarizing noise $\varepsilon=(p/3)(X+Y+Z)$ applied independently to each qubit (Fig.˜5B). Monte Carlo simulation with $10^{5}$ samples per data point (binomial uncertainties $\Delta\varepsilon_{L}/\varepsilon_{L}<5\%$) demonstrates that distance-3 Nash-discovered codes achieve error suppression comparable to surface codes of equal distance, with threshold behavior $\varepsilon_{L}\propto p^{(d+1)/2}$ confirming distance scaling across the sub-threshold regime $p<p_{\mathrm{th}}\approx 1\%$. Surface code baselines at $d=5$ illustrate the additional suppression available at higher distance, with $\varepsilon_{L}\propto p^{3}$ scaling. The rediscovered [[15,7,3]][\![15,7,3]\!] code achieves 47% encoding rate ($k/n=7/15$), substantially higher than distance-3 surface codes encoding one logical qubit per $\sim$9 physical qubits ($k/n\approx 0.11$), providing resource efficiency advantages in overhead-constrained regimes. For hardware-adapted codes with degree $\Delta\leq 3$ on IBM heavy-hex topology [32], the [[15,7,3]][\![15,7,3]\!] code maintains $\varepsilon_{L}<10^{-3}$ at $p=10^{-3}$, meeting fault-tolerance thresholds while fitting existing hardware connectivity—a capability algebraic constructions rarely achieve simultaneously. Computational efficiency enables code discovery for $n\leq 20$ in under 1 second on the hardware described above versus exhaustive search’s $O(2^{n^{2}})$ complexity requiring $10^{13}$ seconds for $n=20$, with the tractable region extending to $n\approx 100$ through optimized implementations.
Figure 5: Metrological Scaling and Error Correction Performance at Hardware-Relevant Scales. (A) Quantum Fisher information $\mathcal{F}_{Q}$ versus number of physical qubits $n$ for Nash-discovered codes at $n=10$, 12, and 15 (filled circles, $\pm 1\sigma$ over $N=5$ independent runs per system size). Solid gray line: Heisenberg limit $\mathcal{F}_{Q}=n^{2}$; dashed gray line: standard quantum limit (SQL) $\mathcal{F}_{Q}=n$. Dash-dot line shows a power-law fit $\mathcal{F}_{Q}\propto n^{1.00}$, confirming that all discovered codes achieve SQL scaling with distance $d=3$. Vertical dash-dot line marks the boundary $n\leq 10$ below which semidefinite programming (SDP) verification remains tractable. (B) Logical error rate $\varepsilon_{L}$ versus physical error rate $p$ under depolarizing noise for Nash-discovered codes at $n=10$ (circles), $n=12$ (squares), and $n=15$ (triangles), all with $d=3$, compared to surface code baselines at $d=3$ (open circles, dashed) and $d=5$ (open diamonds, dashed). Error bars show $\pm 1\sigma$ binomial uncertainty $\sqrt{\varepsilon_{L}(1-\varepsilon_{L})/N}$ with $N=10^{5}$ Monte Carlo samples per data point. Scaling guides $\varepsilon_{L}\propto p^{2}$ ($d=3$) and $\varepsilon_{L}\propto p^{3}$ ($d=5$) are shown as dash-dot lines. Vertical dotted line marks the threshold $p_{\mathrm{th}}\approx 1\%$. Nash-discovered $d=3$ codes achieve error suppression comparable to surface codes of equal distance across the full sub-threshold regime.
Our approach applies game dynamics in a direction opposite to existing quantum game theory: rather than using quantum mechanics to play games [17, 18], we use game theory to construct quantum systems. The “quantum lights out” framework [20] employs single-player optimization metaphors but lacks strategic interactions and equilibrium concepts; reinforcement learning for code discovery [11, 13] explicitly avoids game-theoretic formulation; adversarial noise models [21] analyze robustness but do not construct codes. To our knowledge, no prior work bridges multi-agent game theory and quantum error correction for systematic code discovery through Nash equilibria. Framework extensibility further distinguishes the approach: adding objectives requires only defining new evaluation functions $f_{\text{new}}(G)$—no algorithmic restructuring. Fault-tolerance objectives targeting transversal gates [33], biased noise objectives $f_{\text{bias}}=d_{Z}^{2}(1+d_{X}/d_{Z})$ for asymmetric error channels [34], and hardware topology constraints (heavy-hex, Rydberg arrays, all-to-all connectivity) all enter as modifications to strategy spaces or payoff functions without changing core game logic. This modularity contrasts with algebraic methods requiring distinct mathematical machinery per code family and with reinforcement learning requiring reward and architecture redesign [12]. The transparent graph-to-circuit mapping $G\rightarrow|G\rangle=\prod_{(u,v)\in E}\text{CZ}_{uv}|+\rangle^{\otimes n}$ provides immediate experimental implementation: for the [[15,7,3]][\![15,7,3]\!] code, preparation requires 15 Hadamard gates followed by 21 CZ gates with circuit depth $\propto\Delta(G)=3$.
A natural question is how Nash-discovered codes compare to those found by other computational methods at the same system size. Reinforcement learning approaches [11, 12] have demonstrated code discovery for small systems ($n\leq 10$), but their reported codes do not exceed the parameters discovered here, and they provide no equilibrium certificate or mechanistic explanation of the search trajectory. Random search and genetic algorithms can also discover good codes at small $n$, but lack the convergence guarantees provided by the potential game structure and scale poorly beyond $n\approx 15$ due to the combinatorial explosion of the graph space. The distinctive advantage of the Nash framework is not that it necessarily discovers better codes at fixed nn—at small system sizes, many methods can find good codes—but that it provides (i) a principled convergence criterion via the Nash gap, (ii) mechanistic transparency through equilibrium analysis, and (iii) modular extensibility through objective reconfiguration, together with scalability to $n=100$ where exhaustive and RL approaches become intractable. A systematic head-to-head comparison with RL code search across matched system sizes remains an important direction for future work.
@@
Several limitations warrant acknowledgment. Graph state restriction excludes non-graph-state codes (five-qubit code, Steane code, Bacon-Shor codes). Distance estimation employs heuristics that may underestimate true distance; we validate critical cases with exact weight enumeration when tractable. Approximate Nash equilibria guarantee only local optimality, and the $n=100$ codes require verification against comprehensive code databases to determine novelty. Despite these constraints, the systematic rediscovery of the optimal [[15,7,3]][\![15,7,3]\!] quantum Hamming code validates the methodology, and scalability to $n=100$ demonstrates applicability at hardware-relevant scales. Natural extensions include decoder co-optimization through additional players representing decoder strategies, dynamic objective reconfiguration for degraded hardware, and applications beyond error correction to circuit optimization and resource state design. The interpretable nature of equilibrium-driven discovery may accelerate the transition from theoretical code families to practical implementations by providing mechanistic understanding of why certain topologies succeed—enabling physicists to refine codes based on experimental constraints rather than treating optimization as black-box search.
@@ Acknowledgements.
Appendix A Belief Propagation with Ordered Statistics Decoding for Graph State Stabilizer Codes
We describe the BP+OSD decoder pipeline applied to the graph state stabilizer codes discovered by the Nash equilibrium search. The presentation covers (i) the construction of the parity check matrix from graph state stabilizers, (ii) the belief propagation message schedule, (iii) ordered statistics decoding post-processing, (iv) the Monte Carlo simulation procedure, and (v) distance verification for small codes.
A.1 Stabilizer generators and parity check matrix
A graph state is associated with an undirected graph $G=(V,E)$ on $n=|V|$ vertices. Each vertex $v\in V$ carries a stabilizer generator
$K_{v}=X_{v}\bigotimes_{u\in\mathcal{N}(v)}Z_{u},$ (7)
where $\mathcal{N}(v)$ denotes the neighborhood of $v$ in $G$. The $n$ generators $\{K_{v}\}$ are represented in the binary symplectic formalism by a parity check matrix $H\in\mathbb{F}_{2}^{(n-k)\times 2n}$, partitioned as $H=[H_{X}\mid H_{Z}]$. For a graph state code, the $X$ block encodes which qubits each stabilizer acts on with $X$: because $K_{v}$ applies $X$ only to vertex $v$ itself, $H_{X}=I_{n}$ (the $n\times n$ identity). The $Z$ block encodes the $Z$ support: because $K_{v}$ applies $Z$ to all neighbors $\mathcal{N}(v)$, $H_{Z}=A$, where $A$ is the adjacency matrix of $G$. The parity check matrix therefore takes the explicit form
$H=[I_{n}\mid A]\in\mathbb{F}_{2}^{(n-k)\times 2n}.$ (8)
The number of logical qubits $k=n-\mathrm{rank}_{\mathbb{F}_{2}}(H)$ is determined by the $\mathbb{F}_{2}$-rank of $H$. Logical Pauli operators are constructed from the null space of the symplectic dual of $H$: operators that commute with every row of $H$ but lie outside the row space of $H$ over $\mathbb{F}_{2}$.
A.2 Belief propagation decoding
Syndrome-based decoding maps a measured syndrome $\mathbf{s}\in\mathbb{F}_{2}^{n-k}$ to an estimated error $\hat{\mathbf{e}}$. We apply belief propagation (BP) on the Tanner graph of $H$, treating the $X$ and $Z$ Pauli components independently using the respective rows of $H_{Z}$ and $H_{X}$ as the operative check matrices. Under a depolarizing channel with physical error rate $p$, each qubit independently undergoes an $X$, $Y$, or $Z$ error with probability $p/3$. The channel log-likelihood ratio (LLR) for each variable node $j$ is initialized as
$\lambda_{j}^{(0)}=\log\!\frac{1-q}{q},\quad q=\frac{p}{3}.$ (9)
BP iterates check-to-variable messages
$\mu_{c\to v}^{(t)}=2\,\mathrm{arctanh}\!\left[(-1)^{s_{c}}\prod_{v^{\prime}\in\mathcal{N}(c)\setminus\{v\}}\tanh\!\tfrac{\mu_{v^{\prime}\to c}^{(t-1)}}{2}\right]$ (10)
and variable-to-check messages
$\mu_{v\to c}^{(t)}=\lambda_{j}^{(0)}+\sum_{c^{\prime}\in\mathcal{N}(v)\setminus\{c\}}\mu_{c^{\prime}\to v}^{(t)},$ (11)
where $s_{c}\in\{0,1\}$ is the measured syndrome bit. BP terminates when the estimated syndrome is satisfied or the iteration count reaches 100.
A.3 Ordered statistics decoding post-processing
BP alone frequently fails for codes whose Tanner graphs contain short cycles (trapping sets). Graph state codes derived from locally dense graphs are particularly susceptible because $H_{Z}=A$ can produce many length-4 and length-6 cycles. When BP does not converge, we apply ordered statistics decoding (OSD) as post-processing [31]. OSD ranks all $2n$ variable positions by reliability $|\Lambda_{j}|$ (magnitude of posterior LLR at final BP iteration), row-reduces $H$ to systematic form over $\mathbb{F}_{2}$ in decreasing reliability order, and solves for the remaining bits given the syndrome. At order $w$ (OSD-$w$), we additionally enumerate candidates that flip up to $w$ of the next-most-reliable bits, selecting the minimum-weight candidate. The total complexity is $\mathcal{O}(n^{3}+n^{2}\cdot 2^{w})$. We use OSD-2 throughout this work; at the physical error rates studied ($p\leq 10^{-1}$), OSD-2 post-processing is invoked only for the fraction of trials in which BP does not converge.
A.4 Monte Carlo simulation procedure
We estimate $\varepsilon_{L}$ vs. $p$ by direct Monte Carlo simulation. For each $p\in\{10^{-3},3\times 10^{-3},10^{-2},3\times 10^{-2},10^{-1}\}$:
  1. 1.
    Sample $N$ independent depolarizing error patterns $\mathbf{e}\in\{I,X,Y,Z\}^{\otimes n}$.
  2. 2.
    Compute syndrome $\mathbf{s}=H\mathbf{e}\bmod 2$.
  3. 3.
    Decode with BP+OSD to obtain $\hat{\mathbf{e}}$.
  4. 4.
    Record a logical failure if the residual $\mathbf{e}\oplus\hat{\mathbf{e}}$ commutes with all stabilizer generators but is not in their row space over $\mathbb{F}_{2}$.
The logical error rate estimate is $\hat{\varepsilon}_{L}=N_{\rm fail}/N$, with $\Delta\varepsilon_{L}/\varepsilon_{L}<5\%$ requiring $N\geq 400/\varepsilon_{L}$. For $\varepsilon_{L}\approx 10^{-4}$ we use $N=10^{6}$; otherwise $N=10^{5}$. Error bars in Fig. 3 are $\pm 1\sigma$ of the binomial estimator $\sqrt{\hat{\varepsilon}_{L}(1-\hat{\varepsilon}_{L})/N}$.
A.5 Distance verification for $n\leq 15$ codes
For $n\leq 15$, code distance is computed by exhaustive enumeration: we iterate over all Pauli operators of weight $w=1,2,\ldots$ and check whether each commutes with all stabilizer generators but lies outside the stabilizer group. The minimum such weight is $d$. For the $\llbracket 15,7,3\rrbracket$ code rediscovered by the Nash search, this procedure confirms $d=3$, consistent with Grassl’s code tables [10]. For $n=100$ codes, exhaustive enumeration is intractable; we report lower bounds from the Schlingemann–Werner graph connectivity bound [24], cross-checked by a bounded operator search to depth $w=6$. These lower bounds are estimated to lie within $\pm 1$ of the true distance for representative samples.
Appendix B Objective function coefficient design rationale
Each objective function (Table I) is designed so that ancillary penalty or bonus terms contribute 10–20% of the total reward, keeping SA dynamics driven primarily by the target figure of merit while providing sufficient regularization.
For distance maximization, the cubic exponent $d^{3}$ is motivated by logical error rate scaling $\varepsilon_{L}\propto p^{\lfloor(d+1)/2\rfloor}\sim p^{d}$ near threshold; the steep gradient at high $d$ provides strong selective pressure for distance improvements. The connectivity bonus $\alpha_{\mathrm{conn}}=1.3$ for connected graphs rewards global graph structure associated with larger minimum-weight logical operators. The edge penalty $\beta=0.5$ is calibrated so that, at the typical edge counts found in $n=15$ solutions ($|E|\approx 20$4040), the penalty term contributes roughly 10–15% of the total reward, sufficient to prevent overcrowding without dominating the distance signal.
For hardware-adapted objectives, the reduced exponent $d^{2.5}$ (rather than $d^{3}$) reflects that degree penalties already provide strong regularization; a softer distance gradient allows SA to explore a broader region of graph space before committing to high-degree configurations. The maximum-degree penalty coefficient 5 and average-degree coefficient 2 are set so that, for a typical solution with $\Delta(G)=4$ and $\delta_{\mathrm{avg}}\approx 2.5$, the combined degree penalty is $5\times 4+2\times 2.5=25$ reward units, representing 15–20% of a typical total reward.
For rate-distance, cluster-state, surface-like, and connectivity objectives, analogous design logic applies: penalty and bonus coefficients are set to keep each ancillary term at 10–20% of the primary reward, ensuring that SA dynamics are driven primarily by the objective of interest.